OpenStudy (anonymous):

What is the approximate distance from (4, 1) to (7, 8)?

OpenStudy (saifoo.khan):

use the distance formula!

OpenStudy (anonymous):

\[\sqrt{3^2+7^2}\]

OpenStudy (saifoo.khan):

\[\sqrt{9+49}\]\[\sqrt{58}\]~7.615

OpenStudy (anonymous):

sqr(58)

OpenStudy (anonymous):

...wha?